OpenStudy (anonymous):

Evaluate the following expression for p = -4 and r = 7 -p + 3r + 5(2r + 7) – r + 6p

OpenStudy (anonymous):

\[-p + 3r + 5(2r + 7) – r + 6p\]replace \(p\) by -4, \(r\) by 7 to get \[-(-4)+3\times 7+5(2\times 7+7)-7+6\times (-4)\] then do a bunch of arithmetic

OpenStudy (anonymous):

WHAT